In a vertical circular motion a body of mass 2kg is moving with thread of length 2 meters the change in tension when body reaches lowermost positionnto topmost position is?

Grade:12th pass

1 Answers

Aafiya Shaikh
6 years ago
At top minimum velocity should be such at least centripetal force is equal to the weight of the bodym(v)^2÷r=mgv(at a)=√grSo velocity at top most point(a) is √grAt bottom most point(b) ,by conservation of energy,we get1/2m[v(at b)]^2 = mgh+1/2m[v(at top)^]^2 {here h=2r}So,[v(at b)]^2 = 4gr + grv(at b) = √5grNow,we have,Tension at top most point(a) isT(a) =mv^2/r - mgTension at bottom most point(b) isT(b) = mv^2/r + mgFor change in tension∆T=T(b) - T(a) =m{v(b)}^2/r + mg - m{v(a)^}2/r + mg =5mg + mg - mg + mg =6mg∆T=6×2×10∆T=120 N